OpenStudy (anonymous):

What polynomial has roots of -6, -4, and 1?

OpenStudy (anonymous):

x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

OpenStudy (anonymous):

Since it has the roots -6, -4, and 1 You write it like this (x+6)(x+4)(x-1) Now multiply

OpenStudy (jdoe0001):

\(x=\{-6, -4, 1\}\to \large \begin{cases} x=-6\to {\color{brown}{ x+6}}=0\\ x=-4\to {\color{brown}{ x+4}}=0\\ x=1\to {\color{brown}{ x-1}}=0 \end{cases}\qquad thus \\ \quad \\ \bf ({\color{brown}{ x+6}})({\color{brown}{ x+4}})({\color{brown}{ x-1}})=0\implies ({\color{brown}{ x+6}})({\color{brown}{ x+4}})({\color{brown}{ x-1}})=\textit{original polynomial}\)

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

OpenStudy (anonymous):

so is it B? @Venny @jdoe0001

OpenStudy (anonymous):

Hmm nope! Sorry for the late reply, can you show me how you got b?

OpenStudy (anonymous):

Its ok @Venny its too late.

OpenStudy (anonymous):

Oh, sorry :(

OpenStudy (anonymous):

Its ok!